OpenStudy (anonymous):

A test consists of multiple choice questions worth 2 points and essay questions worth 6 points. The test has 17 questions that are worth a total of 50 points. Use a model to determine how many multiple choice questions are on the test. 10 11 12 13

OpenStudy (anonymous):

2. 2m + 6e = 50 ....(1) m+e=17 ....(2) 2 x (2) : 2m+2e=34 ... (3) (1) - (3 ): 4e=16 e=4, m=13 The answer is 13 :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

your welcome :) if you think i helped you alot please feell free to mark me as your "best response" it will be greatly aprieciated.

OpenStudy (anonymous):

okay i did

OpenStudy (anonymous):

thank you so much :) @alliyahc123

OpenStudy (anonymous):

no problem @gio123